Can you help me solve using the multiplication principle? what is the solution?
-7/8x = -9/10

Answer by linh117(31)

multiply by reciprocal of the coefficient -7/8 on both sides, reciprocal= -8/7
-7/8(-8/7)x = -9/10 (-8/7)
x = -(9)/(10)*(-(8)/(7))
Reduce the expression by removing the common factor of 2 in the denominator of the first term -(9)/(10) and the numerator of the second term -(8)/(7).
-(9)/(5)*-(4)/(7)
Multiply -(9)/(5) by -(4)/(7) to get (36)/(35).
(36)/(35)=x